LHC 2012 - latest news

Week 47: Machine coordinators: Bernhard Holzer, Mike Lamont

Goal of this week: VdM scans & Luminosity production.

Date

Summary

Link to report from morning meeting

 
  • fill 3323 dump at 06:00 Monday morning for start of MD program
08:30 meeting
Sun 25 Nov
  • 01:40 beam dump: spurious BLM trigger
  • 03:30 injection
  • 04:50 stable beams, L0=7.0*10^33
  • 15:50 beams dumped by OP - 168 pb-1
  • Re-fill without issues - fill 3323 - initial lumi 7.1e33 cm-2s-1

9:00 meeting

SPS MKD status (Volker Mertens)

Sat 24 Nov
  • 00:50 End 17th CMS scan
  • 00:55 X/Y scan for ATLAS
  • 02:25 Start VdMs for ALICE
  • 05:00 End ALICE scan
  • 05:10 Dump beams
  • 06:05 Probes in for regular physics
  • 06:15 Steering lines
  • 06:45 No beam from SPS – MKD fault - switch needs to be changed
  • 08:15 Go for a fill (ABT will change switch when LHC in stable beams)
  • 08:45 LHC full…
  • 9:45 Stable beams fill 3318 - phew.
  • 20:20 beam dump by OP, L=147pb^-1
  • 22:16 new injection, 6 bunches swapped between b1 & b2
  • 23:30 stable beams
9:00 meeting
Fri 23 Nov
  • 03:40 loss maps for standard settings
  • SPS problems at MKD3 switch,
  • 07:00 prepare LHC for another vdm scan fill
  • 10:15 Start ramp
  • 10:45 colliding – no problems – clean – good lifetimes
  • 10:55 Stable beams 11:05 beams dumped during ALICE separation - Software interlock (Power converter surveillance). This interlock was not masked as it should have been for the VdM scans....
  • 12:20 ready to inject again but again a problem with the MKD in the SPS
  • 18:00 Injecting physics beam 1
  • 20:00 Start ramp
  • 20:30 Stable beams
  • 20:55 Start 1st CMS VdM scan
08:30 meeting
Thu 22 Nov
  • 3 bunches to flat-top followed by collisions, adjustments etc.
  • 02:00 to 03:00 Collimator set-up
  • Loss maps
  • and back round for full test of settings for Van der Meer run - OK into stable beams without issues
  • 08:15 async. dump test - OK
  • Ramp-down for VdM run but...
  • problem on powercoverter Q10.L8 - access required
  • 14:15 injection for vdm scans
  • 17:30 scnas started in ATLAS, x & y
  • 19:15 CMS solenoid ramp down: cryo problem
  • 23:40 beam dump due to wire scan
08:30 meeting

Wed 21 Nov

  •    waiting for SPS restart
  • investigating of RB QPS disturbance in sector 23
  • 12:35 beam back in SPS…
  • 13:40 SPS MKD switch need to be changed..
  • 19:15 Beam back in SPS
  • 19:15 Cycling after change of active filter on RB.A23
  • 22:00 Pilots in – SPS wrestling with BQM problem - back-out of firmware upgrade required

08:30 meeting

Tue 20 Nov

  • 10:00h validation of new TL colli settings
  • 11:30 IPOC problem, injection not possible
  • 11:45 SPS vacuum leak, magnet has to be exchanged, no beam for 24 h

08:30 meeting

Mon 19 Nov

         03:11 RF problem solved.

         05:38 Stable beams fill 3300. Initial lumi 7.5e33 cm-2s-1.

         14:00 Beam dump by OP, 145 pb-1

         15:00 Access for Wirescanner repair, CMS etc

         18:00 BETS not ok, PLC problem: access needed

         19:00 check of TL collimators: settings are not adequate for Q20 optics in SPS.

         re-calculate correct collimator settings, establish and validate on beam

08:30 meeting